OpenStudy (anonymous):

Find the period ; y=5 cos 1/2 (theta)

OpenStudy (anonymous):

For any Sinusoidal Signal like : \(x(t) = Acos(\omega t + \phi)\) Period is : \[T = \frac{2 \pi}{\omega}\]

OpenStudy (anonymous):

Can you tell what is \(\omega\) value here in your question?

OpenStudy (anonymous):

I really Don't know what's ω !! the question is ; Find the amplitude and period of the equation \[y=5 \cos \frac{ 1 }{2 } \theta\] then graph it # And what i need to know is how to find the period ??

OpenStudy (anonymous):

\(\omega \) is the value attached to your \(\theta\) ..